OpenStudy (anonymous):

HELP , WILL MEDAL which statement is true ? A. -10 + (-9) > 3 + ( -25) B. 7 + ( -7) < -14 + 12 C. 13 + ( -7) > 21 + ( -14 ) D. -18 + 21 = -15 + 12

OpenStudy (anonymous):

which one do you think is right

OpenStudy (anonymous):

you need to sovle all of them to find out do you know how to solve them

OpenStudy (anonymous):

@YoungStudier

OpenStudy (anonymous):

have to simplify both of them ..

OpenStudy (anonymous):

alisa10101 these can be sloved by using a caluclator

OpenStudy (anonymous):

yes you simplfy them

OpenStudy (anonymous):

do you know the less than greater than and equal signs mean

OpenStudy (anonymous):

A and D could be right..right? -10+-9>3+-25 is -19>-22 -18+21=-15+12 is 3=3

OpenStudy (jango_in_dtown):

Add Lhs and Rhs and check the inequalities

OpenStudy (anonymous):

i belive Kobe95 is correct ill give u a meadal

OpenStudy (anonymous):

it gotta be one correct answer

OpenStudy (anonymous):

wait kobe95 check your math for d one yes its the same number but one is positive and one is negitive so not equal

OpenStudy (anonymous):

so A

OpenStudy (anonymous):

yes a is the correct anwser can i get a medal

OpenStudy (anonymous):

sorry didnt notice that

OpenStudy (anonymous):

its ok we all make mistakes